OpenStudy (anonymous):

Solve square root (2x+3) = x + 1 for x 2 +- square root 2 square root 2 4

OpenStudy (souvik):

square both side..

OpenStudy (souvik):

\[\sqrt{2x+3}=x+1\] or \((\sqrt{2x+3})^{2}=(x+1)^{2}\)

ganeshie8 (ganeshie8):

expand right hand side

ganeshie8 (ganeshie8):

\((\sqrt{2x+3})^{2}=(x+1)^{2}\) \(2x+3 =x^2+1 + 2x\) \(2 =x^2\)

OpenStudy (souvik):

or \[x=\pm \sqrt{2}\]

ganeshie8 (ganeshie8):

\(\pm \sqrt{2} = x \) \(x = \pm \sqrt{2} \) its easy.... not as scary as it looks lol, just we need to pen down and keep solving thats all :)